Exercise 18.94 27 of 37 > A Review | Constants Periodic Table Calculate AG at 298 K for the following reactions. Part A Ca(s)Calculate ΔG∘ at 298 K for the following reactions.
Answer #1

Answer:

Lets consider the following reversible reaction: A+B= C+D The following equation relates the standard-state free energy of r

AG = AG° + RT In [C][D] [A][B]

At equilibrium = Det G= 0

At equilibrium condition, ago 49°= -Rt lui [C] ?] TAB TEJ = -KT In .. (als) + (02(g) + + O2(9) CO₂ (s) : Q = [ Ca.(03) - [la